OpenStudy (anonymous):

will give fan and medal 7x/2 < 5. Choose the solution set. {x | x < 10/7} {x | x > 10/7} {x | x < 35/2} {x | x > 35/2}

OpenStudy (anonymous):

help and you will get a medal

OpenStudy (anonymous):

Im thinking

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Sorry no idea

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{7x}{2} < 5 \rightarrow 7x < 2 \times 5 \rightarrow 7x < 10\] \[\rightarrow x <\frac{10}{7}\] Hence option {x | x < 10/7} is the correct answer. @Woody23

OpenStudy (anonymous):

thank you so much @dpasingh

OpenStudy (anonymous):

@Woody23 medal please